# What is the focus of XAI in UI design? Prioritizing UI design principles for enhancing XAI user experience

### v1

@@ 3 XAI User Experience Standards for Non-Expert Users
01 Sajid et al. [11].
Understandability; Satisfaction; Trust; Transparency; Explanation; Trust
02 Samuli et al. [29]
Intelligibility, Comprehensibility, Interpretability; Trust; Transparency; Controllability
03 Markus et al. [30]
Understandable; Satisfaction; Explanation
04 Jasper et al. [10]
Understandable; Persuasion; Correct rate; Accuracy rate
05 Juliana J & Mateus [13]
Adoption rate; Acceptance; Satisfaction; Engagement; Persuasion; Continued use
06 Sule et al. [31]
Usefulness; Naturalness; Trust; Transparency; Controllability
07 Martijn et al. [32]
Effectiveness; Understandability; Trust; Novelty; Satisfaction; Confidence
08 Markus et al. [33]
Trust; Explanation; Satisfaction
09 Robert et al. [34]
Explanation; Satisfaction; Understandability; Curiosity; Trust
10 Nava [35]
Transparency; Scrutability; Trustworthiness; Effectiveness; Persuasiveness; Efficiency; Satisfaction
11 Tim [5]
Coherence; Simplicity; Generality; Truth; Explanation
12 Aniek [36]
Clarity; Parsimony; Completeness; Soundness
13 David & David [16]
Satisfaction; Trust; Predictability; Understandable; Correct rate
14 Nadia & Marco [37]
Trust; Transferability; Causality; Informativeness; Accountability; Transparency
15 Shane et al. [6]
Explanation; Trust; Reliance; Predictability
@@ 3.3 Unique XAI User Experience Level : Understandability, Trust
Satisfaction
This standard measures whether users gain satisfaction during use.
Trust
The standard of trust involves whether users increase their trust in the AI system because of the explanation method.
Persuasiveness
The persuasiveness standard focuses on whether users feel that XAI’s explanation is convincing.
Efficiency
The efficiency standard refers to whether users feel that they have gained higher speed when understanding XAI.
Understandability
The standard of understandability examines whether the content of XAI is easy for users to understand.
@@ 4.1 Design principles for Enhancing XUI User Experience
A: Naturalness
This principle aims to enhance the logic and accuracy of explanations through natural language. It achieves this by using the substantial information content and rapid rationalization characteristic of natural language to generate detailed and logically sound explanations.
B: Responsiveness
The principle of responsiveness aims to dynamically respond to the initial interpretation according to the user’s needs, mainly through progressive disclosures of information to meet the user’s needs. This method not only helps reduce the cognitive load of users but also satisfies users with different depths of understanding.
C: Flexibility
The principle of flexibility encourages the use of multiple different ways of explanation to form a triangular and mutually supporting explanation mechanism and enhance the comprehensiveness and credibility of explanations.
D: Sensitivity
The principle of sensitivity emphasizes the continuous adjustment of explanation principles according to the user’s psychological state and usage scenarios to ensure the adaptability and effectiveness of explanations.
@@ 4.4 Data Collection and Analysis
1
As important as
Means $i$ factors are as important as $j$ factors
3
Slightly more important
Means $i$ factors are slightly more important than $j$ factors
5
Obviously more important
Means $i$ factor is obviously more important than $j$ factor
7
More important
Means $i$ factor is more important than $j$ factor
9
Extremely important
Means $i$ factor is extremely important than $j$ factor
2, 4, 6, 8
Median
The median value of the two adjacent judgments
Count backwards
Relative count backwards
When the $j$ factor is compared with the $i$ factor, the judgment value is $a_{ij}=\frac{1}{a_{ji}}$

### v2

@@ 3 XAI User Experience Standards for Non-Expert Users
01 Sajid et al. [11]. Understandability; Satisfaction; Trust; Transparency; Explanation; Trust
02 Samuli et al. [29] Intelligibility, Comprehensibility, Interpretability; Trust; Transparency; Controllability
03 Markus et al. [30] Understandable; Satisfaction; Explanation
04 Jasper et al. [10] Understandable; Persuasion; Correct rate; Accuracy rate
05 Juliana J & Mateus [13] Adoption rate; Acceptance; Satisfaction; Engagement; Persuasion; Continued use
06 Sule et al. [31] Usefulness; Naturalness; Trust; Transparency; Controllability
07 Martijn et al. [32] Effectiveness; Understandability; Trust; Novelty; Satisfaction; Confidence
08 Markus et al. [33] Trust; Explanation; Satisfaction
09 Robert et al. [34] Explanation; Satisfaction; Understandability; Curiosity; Trust
10 Nava [35] Transparency; Scrutability; Trustworthiness; Effectiveness; Persuasiveness; Efficiency; Satisfaction
11 Tim [5] Coherence; Simplicity; Generality; Truth; Explanation
12 Aniek [36] Clarity; Parsimony; Completeness; Soundness
13 David & David [16] Satisfaction; Trust; Predictability; Understandable; Correct rate
14 Nadia & Marco [37] Trust; Transferability; Causality; Informativeness; Accountability; Transparency
15 Shane et al. [6] Explanation; Trust; Reliance; Predictability
@@ 3.3 Unique XAI User Experience Level : Understandability, Trust
Satisfaction This standard measures whether users gain satisfaction during use.
Trust The standard of trust involves whether users increase their trust in the AI system because of the explanation method.
Persuasiveness The persuasiveness standard focuses on whether users feel that XAI’s explanation is convincing.
Efficiency The efficiency standard refers to whether users feel that they have gained higher speed when understanding XAI.
Understandability The standard of understandability examines whether the content of XAI is easy for users to understand.
@@ 4.1 Design principles for Enhancing XUI User Experience
A: Naturalness This principle aims to enhance the logic and accuracy of explanations through natural language. It achieves this by using the substantial information content and rapid rationalization characteristic of natural language to generate detailed and logically sound explanations.
B: Responsiveness The principle of responsiveness aims to dynamically respond to the initial interpretation according to the user’s needs, mainly through progressive disclosures of information to meet the user’s needs. This method not only helps reduce the cognitive load of users but also satisfies users with different depths of understanding.
C: Flexibility The principle of flexibility encourages the use of multiple different ways of explanation to form a triangular and mutually supporting explanation mechanism and enhance the comprehensiveness and credibility of explanations.
D: Sensitivity The principle of sensitivity emphasizes the continuous adjustment of explanation principles according to the user’s psychological state and usage scenarios to ensure the adaptability and effectiveness of explanations.
@@ 4.4 Data Collection and Analysis
1 As important as Means $i$ factors are as important as $j$ factors
3 Slightly more important Means $i$ factors are slightly more important than $j$ factors
5 Obviously more important Means $i$ factor is obviously more important than $j$ factor
7 More important Means $i$ factor is more important than $j$ factor
9 Extremely important Means $i$ factor is extremely important than $j$ factor
2, 4, 6, 8 Median The median value of the two adjacent judgments
Count backwards Relative count backwards When the $j$ factor is compared with the $i$ factor, the judgment value is $a_{ij}=\frac{1}{a_{ji}}$